# Saturation-Aware Snapshot Compressive Imaging: Theory and Algorithm

### v2

@@ 1 Introduction
Snapshot compressive imaging (SCI) enables fast acquisition of high-dimensional 3D data cubes, such as video [1] and hyperspectral images [2, 3], via specialized optical encoding that maps a 3D data cube into a single 2D snapshot [4]. A critical issue in SCI is sensor saturation, which is more prevalent than in conventional imaging because multiple frames are multiplexed into a single exposure. When the summed intensities exceed the detector’s full-well capacity or analog-to-digital converter (ADC) range, the readout clips (saturates), introducing a nonlinearity that violates the linear SCI model and biases downstream reconstructions, typically causing loss of detail in bright regions and structured artifacts.
Saturation and clipping have been studied in compressed sensing under various nonlinearities, including saturation rejection and consistency methods [5, 6], presaturation errors with noise [7, 8], and quantization effects. However, these analyses apply to generic compressed sensing rather than SCI. In SCI, most prior efforts mitigate saturation empirically, e.g., by designing adaptive or trainable masks [9, 10, 11, 12, 13, 14], incorporating PSF preprocessing to reduce clipping [15]. These works optimize optics or networks but lack recovery guarantees. Despite broad recognition of saturation [16, 17, 15], a theory that explicitly models it within SCI remains absent; existing analyses assume a linear, unsaturated forward model [18, 19]. This gap motivates our development of a theory for SCI under saturation that clarifies how optical encoding (e.g., mask density) and saturation severity jointly govern reconstruction accuracy.
Contributions. We adopt the compression-based SCI recovery framework that is standard in theoretical analyses [18, 19] and develop, to our knowledge, the first theory of SCI under saturation. Specifically: (i) we model clipping as an element-wise nonlinearity and derive a finite-sample recovery bound for compression-based SCI; (ii) the bound makes explicit how reconstruction error depends on mask density and the severity of saturation, yielding a simple mask-design rule; (iii) we instantiate this rule by optimizing Bernoulli mask patterns and pairing them with a new saturation-aware plug-and-play (PnP) reconstruction method; and (iv) experiments on standard video-SCI benchmarks validate the analysis, largely improving robustness and accuracy over conventional masks. Together, these results deepen understanding of saturation in SCI and enable more informed optical and algorithmic design.
@@ 2 Problem statement
Compression-based SCI recovery. We adopt the compression-based recovery framework common in SCI theory [18, 19]. Let $\mathcal{Q}\subset\mathbb{R}^{nB}$ denote a compact signal class. A rate-$r$ compression code $(f,g)$ induces a codebook $\mathcal{C}=\{g(f({\bf x})):{\bf x}\in\mathcal{Q}\}$ with $|\mathcal{C}|\leq 2^{Br}$ and worst-case distortion $\delta=\sup_{{\bf x}\in\mathcal{Q}}\tfrac{1}{nB}\|{\bf x}-g(f({\bf x}))\|_{2}^{2}$. The compression-based estimator (CSP) solves $\hat{{\bf x}}\in\arg\min_{{\bf c}\in\mathcal{C}}\|{\bf y}-{\bf H}{\bf c}\|_{2}^{2}$, as in [18, 19].
@@ 3 Recovery bounds for SCI under saturation
The proof of Theorem 3.1 and its corollaries are provided in Section 4 of the extended version of our paper [20].
In the absence of saturation, we have $p_{s}({\bf x};T)=0$, and Theorem 3.1 reduces to Theorem 1 in [21]. The first term in the bound in (4) reflects the impact of additive noise on the reconstruction error. The second term consists of two components, the latter quantifying the additional reconstruction error due to saturation. As the threshold $T$ increases, the fraction of saturated measurements $p_{s}({\bf x};T)$ and $(\tfrac{B\rho}{2}-T)^{+}\big((\tfrac{B\rho}{2}-T)^{+}+4B\rho\big)$ both decrease. Consequently, the saturation error diminishes with larger $T$, and at $T=\tfrac{B\rho}{2}$ this additional term vanishes entirely.
@@
Corollary 3.2 states that similar to the case of saturation-free measurements, the optimal value of $p^{*}_{T}$ always stays below $0.5$ [19]. Furthermore, it states that as the saturation becomes more severe, to optimize the recovery performance, one needs to lower the density of non-zero entries in the masks.
@@ 4 Proofs
Follow the similar steps in  [19], and $\operatorname{E}[|\mathcal{I}_{s}|]=np_{s}({\bf x};T)$ we have
@@
For the probabilities of $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$, using the union bound together with Hoeffding’s inequality as in [19], we obtain $\operatorname{P}(\mathcal{E}_{1}^{c})+\operatorname{P}(\mathcal{E}_{2}^{c})+\operatorname{P}(\mathcal{E}_{s}^{c})\leq 2^{Br+1}\exp(-n\epsilon_{1}^{2}/(2B^{2}))+\exp(-2n\epsilon_{2}^{2})$.
@@
since $g^{\prime}(p^{*}_{T_{1}};T_{1})=0$, $\Delta_{T_{2}}-\Delta_{T_{1}}<0$, and $\partial p_{s}/\partial p\geq 0$. Because $g^{\prime}(p;T_{2})$ is strictly increasing in $p$ on (0,1/2](0,1/2][19], the unique zero of $g^{\prime}(\cdot;T_{2})$ must occur at some $p^{*}_{T_{2}}>p^{*}_{T_{1}}$. Hence $p^{*}_{T}$ is strictly increasing in $T$, and from (16) we still have $p^{*}_{T}<1/2$ for all $T$, since the right-hand side of (16) is positive.
@@ 5 Experiments
Dataset and Benchmark. We evaluate on six standard grayscale video SCI benchmarks: Kobe, Runner, Drop, Traffic, Aerial, and Crash [22], each at spatial resolution $256\times 256$ with $B=8$. As a baseline, we use the state-of-the-art Plug-and-Play method PnP-FastDVDnet [22].
@@ 5.1 Algorithm
Solving (3) exactly is challenging due to both a non-convex objective and non-convex constraints. In the unsaturated SCI setting, PnP-FastDVDnet [22] provides a robust and efficient approximation of (2) by embedding a pretrained denoiser within an iterative scheme. A key element of this framework is the use of generalized alternating projection (GAP) [23] in place of standard gradient descent. GAP offers two advantages for SCI: (i) it requires no learning-rate tuning (we fix $\mu=1$), and (ii) it efficiently exploits the diagonal structure of the measurement operator $\bf H$}.
@@ 5.3 SCI recovery under measurements saturation
Figure 3: PSNR of $\|{\bf x}-{\bf\hat{x}}\|$ under different masks with elements that are i.i.d.  $\mathrm{Bern}(p)$. Solid lines represent results obtained with the proposed SAPnet, while dotted lines correspond to PnP-FastDVDnet. Blue curves indicate a saturation threshold of $T/B=0.25$, and Orange curves indicate $T/B=0.5$.
@@
For the setup with small measurement noise, our theoretical analysis in Section 3.1 demonstrates that the results remain robust under additive perturbations. In Fig. 4, we simulate noisy measurements with noise level $\sigma=10$ and compare the baseline method with the proposed SAPnet. The results show that SAPnet maintains robustness against noise, consistently achieving higher performance with optimized masks. While the optimal mask probability grows slightly compared to the ideal (noise-free) case, this behavior is consistent with the detailed explanation provided in Section 3 of [24], and corresponds to the unsaturated noisy measurement results.

### v3

@@ 1 Introduction
Snapshot compressive imaging (SCI) enables fast acquisition of high-dimensional 3D data cubes, such as video [llull2013coded] and hyperspectral images [gehm2007single, Wagadarikar08CASSI], via specialized optical encoding that maps a 3D data cube into a single 2D snapshot [yuan2021snapshot]. A critical issue in SCI is sensor saturation, which is more prevalent than in conventional imaging because multiple frames are multiplexed into a single exposure. When the summed intensities exceed the detector’s full-well capacity or analog-to-digital converter (ADC) range, the readout clips (saturates), introducing a nonlinearity that violates the linear SCI model and biases downstream reconstructions, typically causing loss of detail in bright regions and structured artifacts.
Saturation and clipping have been studied in compressed sensing under various nonlinearities, including saturation rejection and consistency methods [laska2011, rencker2019sparse], presaturation errors with noise [foucart2018sparse, banerjee2024likelihood], and quantization effects. However, these analyses apply to generic compressed sensing rather than SCI. In SCI, most prior efforts mitigate saturation empirically, e.g., by designing adaptive or trainable masks [Wang_2023_ICCV, Zhang22optics, Metzler_2020_CVPR, martelneuralsensor, CAI2025113913, zhao2024sasa], incorporating PSF preprocessing to reduce clipping [zhang2024anti]. These works optimize optics or networks but lack recovery guarantees. Despite broad recognition of saturation [correa2016spatiotemporal, diaz2019adaptive, zhang2024anti], a theory that explicitly models it within SCI remains absent; existing analyses assume a linear, unsaturated forward model [jalali2019snapshot, zhao2025theoretical]. This gap motivates our development of a theory for SCI under saturation that clarifies how optical encoding (e.g., mask density) and saturation severity jointly govern reconstruction accuracy.
Contributions. We adopt the compression-based SCI recovery framework that is standard in theoretical analyses [jalali2019snapshot, zhao2025theoretical] and develop, to our knowledge, the first theory of SCI under saturation. Specifically: (i) we model clipping as an element-wise nonlinearity and derive a finite-sample recovery bound for compression-based SCI; (ii) the bound makes explicit how reconstruction error depends on mask density and the severity of saturation, yielding a simple mask-design rule; (iii) we instantiate this rule by optimizing Bernoulli mask patterns and pairing them with a new saturation-aware plug-and-play (PnP) reconstruction method; and (iv) experiments on standard video-SCI benchmarks validate the analysis, largely improving robustness and accuracy over conventional masks. Together, these results deepen understanding of saturation in SCI and enable more informed optical and algorithmic design.
@@ 2 Problem statement
Compression-based SCI recovery. We adopt the compression-based recovery framework common in SCI theory [jalali2019snapshot, zhao2025theoretical]. Let $\mathcal{Q}\subset\mathbb{R}^{nB}$ denote a compact signal class. A rate-$r$ compression code $(f,g)$ induces a codebook $\mathcal{C}=\{g(f({\bf x})):{\bf x}\in\mathcal{Q}\}$ with $|\mathcal{C}|\leq 2^{Br}$ and worst-case distortion $\delta=\sup_{{\bf x}\in\mathcal{Q}}\tfrac{1}{nB}\|{\bf x}-g(f({\bf x}))\|_{2}^{2}$. The compression-based estimator (CSP) solves $\hat{{\bf x}}\in\arg\min_{{\bf c}\in\mathcal{C}}\|{\bf y}-{\bf H}{\bf c}\|_{2}^{2}$, as in [jalali2019snapshot, zhao2025theoretical].
@@ 3 Recovery bounds for SCI under saturation
The proof of Theorem 3.1 and its corollaries are provided in Section 4 of the extended version of our paper [zhao2025saturation].
In the absence of saturation, we have $p_{s}({\bf x};T)=0$, and Theorem 3.1 reduces to Theorem 1 in [zhao2023theoretical]. The first term in the bound in (4) reflects the impact of additive noise on the reconstruction error. The second term consists of two components, the latter quantifying the additional reconstruction error due to saturation. As the threshold $T$ increases, the fraction of saturated measurements $p_{s}({\bf x};T)$ and $(\tfrac{B\rho}{2}-T)^{+}\big((\tfrac{B\rho}{2}-T)^{+}+4B\rho\big)$ both decrease. Consequently, the saturation error diminishes with larger $T$, and at $T=\tfrac{B\rho}{2}$ this additional term vanishes entirely.
@@
Corollary 3.2 states that similar to the case of saturation-free measurements, the optimal value of $p^{*}_{T}$ always stays below $0.5$ [zhao2025theoretical]. Furthermore, it states that as the saturation becomes more severe, to optimize the recovery performance, one needs to lower the density of non-zero entries in the masks.
@@ 4 Proofs
Follow the similar steps in  [zhao2025theoretical], and $\operatorname{E}[|\mathcal{I}_{s}|]=np_{s}({\bf x};T)$ we have
@@
For the probabilities of $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$, using the union bound together with Hoeffding’s inequality as in [zhao2025theoretical], we obtain $\operatorname{P}(\mathcal{E}_{1}^{c})+\operatorname{P}(\mathcal{E}_{2}^{c})+\operatorname{P}(\mathcal{E}_{s}^{c})\leq 2^{Br+1}\exp(-n\epsilon_{1}^{2}/(2B^{2}))+\exp(-2n\epsilon_{2}^{2})$.
@@
since $g^{\prime}(p^{*}_{T_{1}};T_{1})=0$, $\Delta_{T_{2}}-\Delta_{T_{1}}<0$, and $\partial p_{s}/\partial p\geq 0$. Because $g^{\prime}(p;T_{2})$ is strictly increasing in $p$ on (0,1/2](0,1/2][zhao2025theoretical], the unique zero of $g^{\prime}(\cdot;T_{2})$ must occur at some $p^{*}_{T_{2}}>p^{*}_{T_{1}}$. Hence $p^{*}_{T}$ is strictly increasing in $T$, and from (16) we still have $p^{*}_{T}<1/2$ for all $T$, since the right-hand side of (16) is positive.
@@ 5 Experiments
Dataset and Benchmark. We evaluate on six standard grayscale video SCI benchmarks: Kobe, Runner, Drop, Traffic, Aerial, and Crash [PnP_fastdvd], each at spatial resolution $256\times 256$ with $B=8$. As a baseline, we use the state-of-the-art Plug-and-Play method PnP-FastDVDnet [PnP_fastdvd] and model-based method GAP-TV [yuan2016generalized].
@@ 5.1 Algorithm
Solving (3) exactly is challenging due to both a non-convex objective and non-convex constraints. In the unsaturated SCI setting, PnP-FastDVDnet [PnP_fastdvd] provides a robust and efficient approximation of (2) by embedding a pretrained denoiser within an iterative scheme. A key element of this framework is the use of generalized alternating projection (GAP) [liao2014generalized] in place of standard gradient descent. GAP offers two advantages for SCI: (i) it requires no learning-rate tuning (we fix $\mu=1$), and (ii) it efficiently exploits the diagonal structure of the measurement operator $\bf H$}.
@@ 5.2 Saturated measurements’ distribution
Table 2: PSNR (dB) comparison among GAP-TV [yuan2016generalized] (model-based), PnP-FastDVDnet [PnP_fastdvd] (pretrained deep denoiser, represent by PnP), and SAPnet (ours) on six grayscale benchmarks under different saturation levels $T/B$ with optimized mask. Bold indicates the best performance and underline indicates the second-best.
Dataset $T/B=0.25$ $T/B=0.5$ $T/B=0.75$
GAP-TV PnP SAPnet GAP-TV PnP SAPnet GAP-TV PnP SAPnet
Runner 25.57 31.88 37.04 27.39 37.01 37.90 27.80 37.92 37.94
Drop 24.76 12.37 42.26 29.79 36.29 43.26 31.57 43.03 43.24
Crash 22.25 20.60 26.13 23.98 26.25 26.86 24.29 26.87 26.86
Traffic 18.45 22.73 26.75 19.35 27.31 27.92 19.43 27.98 27.99
Kobe 21.95 28.94 31.70 22.28 32.58 32.75 22.29 32.80 32.79
Aerial 24.33 24.56 26.73 25.42 27.83 27.91 25.54 27.90 28.13
@@ 5.3 SCI recovery under measurements saturation
Figure 3: PSNR of $\|{\bf x}-{\bf\hat{x}}\|$ under Bernoulli masks with i.i.d. entries $\mathrm{Bern}(p)$. Solid lines correspond to the proposed SAPnet, dashed lines to GAP-TV, and dotted lines to PnP-FastDVDnet. Blue curves show results for saturation level $T/B=0.25$, while orange curves show results for $T/B=0.5$.
@@
For comparison, we also report results of GAP-TV, a classical model-based baseline that does not explicitly model measurement saturation. As shown in Fig. 3, GAP-TV follows the same overall trend as SAPnet across datasets and mask densities, with reconstruction quality improving as saturation weakens and the mask probability approaches its optimal value, which is consistently below $0.5$. As saturation increases, the optimal mask probability further decreases. However, GAP-TV consistently achieves lower PSNR than SAPnet at all saturation levels, with a larger gap in strongly saturated cases. This observation is consistent with the performance results in Table 2, and shows the benefit of explicitly enforcing saturation consistency rather than relying only on a total-variation prior.
@@
For the setup with small measurement noise, our theoretical analysis in Section 3.1 demonstrates that the results remain robust under additive perturbations. In Fig. 4, we simulate noisy measurements with noise level $\sigma=10$ and compare the baseline method with the proposed SAPnet. The results show that SAPnet maintains robustness against noise, consistently achieving higher performance with optimized masks. While the optimal mask probability grows slightly compared to the ideal (noise-free) case, this behavior is consistent with the detailed explanation provided in Section 3 of [zhao2024untrained], and corresponds to the unsaturated noisy measurement results.